OpenStudy (anonymous):

d/dx [(sin(pi/2) + cos(pi/2)]^5

zepdrix (zepdrix):

\[\Large\rm \left[\sin(\pi/2)+\cos(\pi/2)\right]^5=\text{#}\]\[\Large\rm \frac{d}{dx}\text{#}=0\]

zepdrix (zepdrix):

There is no x variable present. Did you want to take derivative and then evaluate it at x=pi/2 instead maybe?

OpenStudy (anonymous):

so its 0 ?

zepdrix (zepdrix):

Yes c: There is no x variable, the entire thing is constant with respect to x. Derivative, or rate change, of a constant is zero.

OpenStudy (anonymous):

thanks!